### v1

@@ 3 Data and Methodology
3.1 Framework
Based on the reviewed literature, the goal of this study is:
  1. 1.
    Assess the ability of selected deep learning models to accurately reproduce streamflow estimates for the given catchments.
  2. 2.
    Examine several methods of introducing static catchment attributes as additional inputs, such as the Batch-Static and Stacked Ensemble Architecture.
  3. 3.
    Compare selected deep learning model and data configuration to evaluate model strengths and failures when attempting to predict extreme rainfall events such as floods.
  4. 4.
    Address the failure of conventional models by introducing a novel conditional ensemble deep learning model that employs features of quantile regression.
Figure 2 presents the framework, which features three key stages. Using our framework, we examine deep learning models with catchment data strategies as given below:
  1. 1.
    The Individual strategy features a separate model for each base station, as shown in Figure 4.
  2. 2.
    The Batch-Indicator strategy includes base stations from the entire region, where data from all the basins are concatenated for a single model as shown in Figure 5.
  3. 3.
    Batch-Static strategy features data from the entire region with static input, where each input vector features static variables as shown in Figure 6.
  4. 4.
    The Ensemble catchment strategy (Figure 7) combines forecasts from two models: one capturing temporal sequences and the other static features. These forecasts are concatenated and processed through a regression layer to generate final predictions. We investigate if the overall prediction accuracy improves by leveraging the strengths of both temporal and static feature models, enhancing the model’s ability to make more accurate and robust predictions.
In the initial stage, we assess the performance of four distinct model architectures: i.e. an Individual architecture, a Batch Indicator, a Batch Static, and an Ensemble model. Then we advance to the second stage, where we evaluate the performance of four deep learning models within the chosen architecture, based on the summary provided in 2.2: LSTM, CNN, BD-LSTM, and ED-LSTM. In the third stage, we utilise the conditional ensemble model that incorporates features from quantile regression. We compare it against the previously identified best-performing models to derive valuable insights about predictive capabilities and reliability.
@@ 3.2 CAMELS dataset
We make use of the catchment attributes and meteorological (CAMELS) dataset [78] to demonstrate the use of deep learning models in flood predictions. CAMELS is a set of data containing hydrometeorological time series and landscape attributes for 222 catchments in Australia, Figure 1 presents an overview of the approximate locations.
@@
Table 1: Summary of available data in CAMELS dataset
@@
Summary Mean streamflow mm $d^{-1}$
Summary Sensitivity of streamflow to rainfall -
Summary Ratio of streamflow to precipitation -
Summary Frequency of high flow days d y-1
Summary Average duration of high flow days
Summary Frequency of low flow days d y-1
Summary Frequency of days with zero flow d y-1
The CAMELS dataset is suitable for “large-sample hydrology” [79, 80] and for evaluation of deep learning models [81]. The CAMELS time series data consists of streamflow and 18 climatic variables at daily timesteps. The start of the time series is the 1st of January, 1911; however, since more stations were added over time, the starting date of each time series per station is different. For example, for Station A, streamflow data may range from 1 January 1950 to 1 January 2020 , but for Station B streamflow data may range from 1 January 1980 to 1 January 2020. As a result, if the analysis is across stations , we must take the dates that are all in common.
The landscape attribute gives data and descriptions of attributes related to geology, soil, topography, land cover, anthropogenic influence and hydroclimatology for all 222 catchments. This is static information, so each catchment will only have one value for each attribute. The metadata is also available, which describes the latitude, longitude, boundaries of each catchment. We provide a summary of the different data types in Table 1.
The dataset contains a variety of data that may not be relevant to the problem of flood prediction, so we select only the relevant time series and static variables. In all our models, the input time series includes i.) daily cumulative precipitation, ii.) daily minimum air temperature, iii.) daily maximum air temperature, iv.) the time of year as sine/cosine functions.
Furthermore, we selected certain catchment characteristics for the models that require static data. These include i.) mean streamflow, ii.) sensitivity of streamflow to rainfall. iii.) ratio of streamflow to precipitation. iv.) frequency of high flow days. v.) average duration of high flow. vi.) frequency of low flow days. vii.) frequency of days with zero flow. These catchment attributes include climatic and vegetation indices, as well as soil and topographical properties.
@@ 3.3 Data Reconstruction
In order to use deep learning models, the original time-series data (which is common in climate and hydrological science) needs to be transformed (embedded) to be used for either single-step or multistep-ahead prediction. For this, the data needs to be processed using a sliding window approach, and the size of the window is crucial for this step. Taken’s theorem demonstrates that any smooth coordinate transformation does preserve the important properties of the original time-series data [82] with an appropriate embedding dimension. Thus, if the original univariate time-series data is of the form:
$Y=[x(1),x(2),...,x(t),...,x(N)]$ (3)
where, we observe $x(t)$ at time $t$. We can create an  textitembedded phase space via sliding windows:
@@
where, $T$ is the time delay, $D$ is the embedding dimension (window-size); this is valid for all values of $t$ from $t=0,1,2,...,N-DT-1$ and $N$ is the length of the data. Assuming the original attractor is of dimension $d$, then $D=2d+1$ would be sufficient to capture the important characteristics of the original data in the embedded phase space [82]. We will use the embedded phase space for training the deep learning models where $D$ will be used to define the input size (neurons) of the respective deep learning models. Essentially, a univariate time series vector gets transformed into a matrix that is used for training and testing the model. Rather than using the entire univariate time series, the embedded phase space enables a sliding window approach for training the respective deep learning model, where the input is defined by the size of the sliding window. In the case of a multivariate time series, the same approach can be used, as shown in Figure 3.
@@
We need to transform the two-dimensional dataset of features and temporal sequences (daily time interval) into a three-dimensional format through a sliding window technique. A window of size of $N$ slides over the 2-dimensional dataset given by $6\times T$; where 6 is the number of input variables, and $T$ is the length of the dataset (number of days for available data). This creates a 3D dataset of $M\times N\times 6$; where $M=T-N+1$, i.e. the number of times the window can slide over the dataset. As a result, we create a feature input dataset $X_{i}=[x_{1},x_{2},...,x_{5}][s_{1}],\;i=1...N$ and target (prediction) $Y_{j}=[y_{1},y_{2},...,y_{N_{o}ut}],\;j=1...N-out$, as shown in Figure 3.
Figure 2: Framework for comparing the different models for the different architectures.
@@
Figure 3: Embedding: transformation of a two-dimensional temporal dataset of features (daily time interval) into a three-dimensional format through a sliding window technique.
Figure 4: Individual model per station. [x1, x2, x3, x4] refers to the collection of time series variables at a point in time. [y1, y2, y3, y4] are the corresponding prediction horizons (multi-step prediction). …
Figure 5: Batch model. [x1, x2, x3, x4] refers to the collection of time series variables at a point in time. [y1, y2, y3, y4] are the corresponding prediction horizons (multi-step prediction).
Figure 6: Batch model with static data. [x1, x2, x3, x4] refers to the collection of time series variables at a point in time. The arrangement of [s1] is for models requiring static features as a one-time input. [y1, y2, y3, y4] are the corresponding prediction horizons (multi-step prediction).
@@
Figure 7: Stacked Ensemble strategy combining Temporal and Static models using Linear Regression.
3.4 Quantile-LSTM: conditional ensemble deep learning model
@@
Our conditional ensemble deep learning model (Quantile-LSTM) capitalises on quantile regression using a 1D-LSTM model (Figure 8). We will use the same data as for the deep learning models for prediction of the flow-duration exceedance probability ($\alpha$), which is the cumulative frequency curve that shows the probability of streamflow exceeding a given threshold for the data period [83]. In essence, this initial model predicts the probability that a flood of a certain magnitude occurs in the prediction window.
Quantile-LSTM is based on the quantiles of a distribution, and will offer a comprehensive understanding of prediction uncertainty. We strategically use three separate Quantile-LSTM models to optimize the performance for different flood probability scenarios. Depending on the predicted probability, the LSTM models provide three separate n-step-ahead predictions; i.e. we train a Quantile-LSTM at a quantile level ($q=0.95$), when the predicted flood exceedance probability is above a threshold ($\alpha>0.95$), ensuring accurate forecasting in high-probability flood events. Similarly, if $\alpha$ falls between 0.7 and 0.95, we use a Quantile-LSTM with $q=0.70$. In contrast, when $\alpha<0.70$, we train a standard LSTM model since this region does not feature extreme values in data. This tripartite approach allows the model to adapt to varying flood risk levels, capitalising on the flexibility and specificity of the Quantile-LSTM in capturing different aspects of the flow-duration curve. Figure 8 presents the details of the switching mechanism for the conditional ensemble model which decides which Quantile-LSTM model to be used. We will evaluate this approach against the observed data and the best-performing deep learning models.
Figure 8: Conditional ensemble model utilising Quantile-LSTM.
@@ 3.5 Experiment Setup
In this section, we provide technical details for the baseline deep learning models (LSTM, BD-LSTM, ED-LSTM, and CNN) and our proposed models that include Stacked Ensemble and Quantile-LSTM. We determined model hyperparameters such as hidden neurons, learning rate and epochs from trial experiments, as shown in Table 2. We note that all the tested models implement a multi-step time series prediction with 5 inputs and 5 outputs (time steps). We train the model with a maximum time of 150 epochs using an Adam optimiser [84] with default hyperparameters, such as the initial learning rate.
In order to develop regional-based models, we separate the CAMELS data based on the state. All time-series data are scaled using a min-max-scaler, i.e. scaled in the range [0,1]. In each basin, there are about 10,000 data points representing daily data spanning from 1990 to 2015. We use 60% for training and 40% for testing. We used the first 60 % of the data from 1990 as training data and the remaining for testing. We note that although there may be a climate bias in taking a block of data spanning more than a decade, this approach ensures reproducibility in results when compared to shuffling the data.
It is necessary to use a variety of performance metrics for model benchmarking, as no one evaluation metric can fully capture the consistency, reliability, accuracy, and precision of a stream flow model. The Nash–Sutcliffe-Efficiency (NSE) is the hydrology version of coefficient of determination, similar to $R^{2}$, which can be less than zero. Silva et al. [85] have suggested the Squared Error Relevance (SER) - (SER-$\alpha$%) for extreme events. The SER calculates the Root Mean Squared Error (RMSE) of test windows that contain the $\alpha$ percentile of the data, measuring the ability of a model to predict rare data points. For example, SER-1% would find the RMSE of windows that contain the 1% most extreme stream flows for a given basin. In summary, we use the following metrics to evaluate prediction performance accuracy over multiple-steps (prediction horizon).
@@
Table 2: Model Configuration
Models Hidden Neurons Comments
LSTM 20 Cells Single Layered
BD-LSTM 20 Cells Single Layered
ED-LSTM 20 Cells Single Layered
CNN 64 Cells Single Layered
Quantile-LSTMq=0.5 20 cells Conditional Ensemble Model
Quantile-LSTMq=0.7 20 cells Conditional ensemble Model
LSTM 20 cells Conditional Ensemble Model
Table 3: Comparison of catchment strategies using LSTM model for South Australia
@@
Table 4: Architecture wise- Model Comparisons to track best performing model (South Australia)
@@
Quantile-LSTM 1.6509 1.0572 0.5826 0.3604 0.1963 0.1220 0.1006 0.0929
±0.0743 ±0.0418 ±0.0181 ±0.0093 ±0.0040 ±0.0021 ±0.0014 ±0.0011
Table 5: Quantile-LSTM results for major states in Australia
State Number of Stations SER 1% SER 2% SER 5% SER 10% SER 25% SER 50% SER 75% RMSE
SA 9 1.6509 1.0572 0.5826 0.3604 0.1963 0.1220 0.1006 0.0929
±0.0743 ±0.0418 ±0.0181 ±0.0093 ±0.0040 ±0.0021 ±0.0014 ±0.0011
NSW 33 1.4779 0.9425 0.4808 0.2770 0.1285 0.0695 0.0490 0.0401
±0.2421 ±0.1556 ±0.0805 ±0.0459 ±0.0227 ±0.0164 ±0.0134 ±0.0121
QLD 61 1.4556 0.9937 0.5355 0.3117 0.1396 0.0727 0.0506 0.0425
±0.2408 ±0.1519 ±0.0961 ±0.0678 ±0.0476 ±0.0263 ±0.0214 ±0.0122
VIC 72 1.0603 0.6685 0.3685 0.2241 0.1165 0.0693 0.0528 0.0451
±0.2817 ±0.1823 ±0.1041 ±0.0704 ±0.0502 ±0.0286 ±0.0230 ±0.0134
WA 16 2.3148 1.7796 1.1950 0.8078 0.4384 0.2735 0.2132 0.1891
±0.1550 ±0.0827 ±0.0402 ±0.0331 ±0.0193 ±0.0133 ±0.0118 ±0.0113
NT 16 1.7982 1.2585 0.7505 0.4838 0.2431 0.1509 0.1241 0.1104
±0.0251 ±0.0183 ±0.0199 ±0.0168 ±0.0121 ±0.0100 ±0.0094 ±0.0092
TAS 12 1.4647 1.0495 0.6412 0.4272 0.2419 0.1526 0.1129 0.0876
±0.0791 ±0.0591 ±0.0379 ±0.0243 ±0.0145 ±0.0119 ±0.0092 ±0.0081
ACT 3 1.7321 1.0715 0.4775 0.2528 0.1147 0.0607 0.0413 0.0317
±0.0564 ±0.0288 ±0.0158 ±0.0085 ±0.0035 ±0.0018 ±0.0016 ±0.0016

### v2

@@ 3.1 CAMELS dataset
We use the CAMELS dataset [41] to demonstrate the use of deep learning models in streamflow prediction. CAMELS data contains hydrometeorological time series and landscape attributes for 222 catchments in Australia, Figure 1 presents an overview of the approximate locations.
@@
Table 1: Time series and static (landscape attribute) in the CAMELS dataset selected for our study.
@@
Static Mean streamflow mm $d^{-1}$
Static Sensitivity of streamflow to rainfall -
Static Ratio of streamflow to precipitation -
Static Frequency of high flow days d y-1
Static Average duration of high flow days
Static Frequency of low flow days d y-1
Static Frequency of days with zero flow d y-1
The CAMELS dataset is suitable for “large-sample hydrology” [83, 84] featuring time series of streamflow and 18 climatic variables at daily time steps. The time series data was recorded from January 1911 and the addition of more stations over time featured a new beginning date per station. For example, for Station A, streamflow data would range from 1 January 1950 to 1 January 2020, but for Station B streamflow data would range from 1 January 1980 to 1 January 2020. Therefore, since our analysis is across stations, we must make a snapshot of the timeline with data present, and hence we selected data from 1st January 1980 January to 1st January 2020.
The landscape attribute gives data and descriptions of attributes related to geology, soil, topography, land cover, anthropogenic influence and hydroclimatology for all 222 catchments. This is considered static information, where each catchment features only one value for each attribute. The metadata describes the latitude, longitude, and boundaries of each catchment. We provide a summary of the different attribute types in Table 1.
The dataset contains a variety of data that may not be relevant to the problem of flood prediction, so we select only the relevant time series data and static variables. In all our models, the input time series includes i.) daily cumulative precipitation, ii.) daily minimum air temperature, iii.) daily maximum air temperature, iv.) the time of year as sine/cosine functions as shown in Table 1. Furthermore, we selected certain catchment characteristics for the models that require static data. These include i.) mean streamflow, ii.) sensitivity of streamflow to rainfall. iii.) ratio of streamflow to precipitation. iv.) frequency of high flow days. v.) average duration of high flow. vi.) frequency of low flow days. vii.) frequency of days with zero flow as given in Table 1. These catchment attributes include climatic and vegetation indices, as well as soil and topographical properties.
@@ 3.2 Data Reconstruction
In order to use deep learning models, the original time-series data needs to be transformed (embedded) for both single-step and multistep-ahead prediction strategies. Therefore, the data needs to be processed using a sliding window approach, and the size of the window is crucial for this step. Taken’s theorem demonstrates that any smooth coordinate transformation does preserve the important properties of the original time-series data [85] with an appropriate embedding dimension. Thus, if the original univariate time-series data is of the form:
$Y=[x_{1},x_{2},...,x_{t},...,x_{N}]$ (3)
where, we observe $x_{t}$ at time $t$. We can create an embedded phase space via sliding windows:
@@
where, $T$ is the time delay, $D$ is the embedding dimension (window-size); this is valid for all values of $t$ from $t=0,1,2,...,N-DT-1$ and $N$ is the length of the data. Assuming the original attractor is of dimension $d$, then $D=2d+1$ would be sufficient to capture the important characteristics of the original data in the embedded phase space [85]. We use the embedded phase space to train the deep learning models, where $D$ represents the models’ input size (neurons). Essentially, a univariate time series vector gets transformed into a matrix that is used for training and testing the model. Rather than using the entire univariate time series, the embedded phase space enables a sliding window approach for training the respective deep learning model, where the size of the sliding window defines the input. In the case of a multivariate time series, the same approach can be used, as shown in Figure 3.
@@
We need to transform the two-dimensional dataset of features and temporal sequences (daily time interval) into a three-dimensional format through a sliding window technique, as shown in Figure 3. We use a sliding window (size of $D\times F$) sliding over the 2-dimensional dataset given by $F\times D$; where $F$ is the number of multivariate features. $N$ is the total length of the dataset, given by the number of days.
Figure 2: Framework for comparing the different models for the different architectures (Stage 2 and 3) after acquiring and processing data (Stage 1). Stage 4 features our ensemble quantile deep learning model that predicts 2 extreme quantiles ($5^{th}$ and $95^{th}$ quantiles) along with regular streamflow prediction. We then utilise the model from Stage 4 to selected catchments across Australian states in Stage 5 and show results using the respective metrics (RMSE, NSE and SER). Finally, in Stage 6, we provide a hydrograph visualisation of flood probability and streamflow prediction for selected catchments.
@@
Figure 3: Embedding: the transformation of a two-dimensional temporal dataset of features (daily time interval) into a three-dimensional format through a sliding window technique (windowing).
(a) Individual models trained for each (individual) station separately.
(b) Batch-temporal: Single model for a subset of stations using temporal features only.
(c) Batch-static: a single model for a subset of stations using both temporal and static features.
@@
(d) Stacked Ensemble strategy: Combining Temporal and Static models using a dense neural network.
Figure 4: Configurations
3.3 Framework
The contribution of our study is through fulfilling the following goals in order to accurately reproduce streamflow estimates for the given catchments.
  1. 1.
    Incorporate landscape attributes about the catchment (static input) as an additional feature with the time series data for the respective deep learning and ensemble learning models.
  2. 2.
    Evaluate selected deep learning model and data configuration for streamflow prediction across selected Australian catchments.
  3. 3.
    Develop ensemble quantile-based deep learning model for uncertainty quantification in streamflow prediction.
  4. 4.
    Utilise streamflow prediction to create flood probability based on flood frequency analysis.
  5. 5.
    Compare flood probability based on the historical flooding events of selected catchments for risk analysis.
Figure 2 presents the framework, which features deep learning models with catchment data strategies as given below:
  1. 1.
    The Individual strategy features a separate model for each base station, as shown in Figure 4(a).
  2. 2.
    The Batch-Temporal strategy includes base stations from the entire region, where data from all the basins are concatenated for a single model as shown in Figure 4(b).
  3. 3.
    Batch-Static strategy features data from the entire region with static input, where each static input features landscape attributes as shown in Figure 4(c).
  4. 4.
    The Ensemble catchment strategy (Figure 4(d)) combines forecasts from two models: one capturing temporal sequences and the other static features. These forecasts are concatenated and processed through a regression layer to generate final predictions. We investigate if the overall prediction accuracy improves by leveraging the strengths of both temporal and static feature models, enhancing the model’s ability to make more accurate and robust predictions.
We begin with data processing that includes time series data reconstruction for use in deep learning models, as shown in Stage 1 of Figure 2. We implement data reconstruction (windowing) and process precipitation-deficit, temperature, and antecedent streamflow as temporal data. The static data includes catchment streamflow properties such as frequency and duration of high/low streamflow and rainfall sensitivity. In Stage 2, we assess the performance of four distinct catchment strategies: i.e. an Individual architecture, Batch-Temporal, Batch-Static, and Stacked-Ensemble model using an LSTM model. In Stage 3, we evaluate the performance of four deep learning models including LSTM, CNN, bidirectional LSTM (BD-LSTM), and encoder-decoder LSTM (ED-LSTM).
In Stage 4, we train the ensemble model, where we implement quantile regression in the output layer of the deep learning model (Figure 5). The ensemble features two Quantile-LSTM and a regular LSTM model to quantify the uncertainty in data via a confidence interval over the prediction horizon where the upper and lower bound is predicted by the Quantile-LSTM models, along with the conventional LSTM model for regular prediction.
In Stage 5, we compare the respective models and apply the selected catchments across the Australian states including NSW, Queensland, and Victoria. In Stage 5, we show results using the respective metrics. Finally, in Stage 6, we provide a hydrograph visualisation of flood probability and streamflow prediction for selected catchments.
3.4 Quantile-LSTM: ensemble deep learning model
@@
Our ensemble deep learning model capitalises on quantile regression using an LSTM model known as Quantile-LSTM (Figure 5). The Quantile-LSTM is based on the quantiles [86] of a distribution, which can enable a comprehensive understanding of prediction uncertainty when compared to the conventional LSTM model. We strategically use two separate Quantile-LSTM models to optimise for upper quantile $(q=0.95)$ and lower quantile $(q=0.05)$ of the $90\%$ confidence interval. The confidence interval quantifies the data uncertainty in the streamflow prediction over the n-step prediction horizon. In addition, we train a regular-LSTM model using the mean-squared regression for predicting the expected value of streamflow over the prediction horizon. In order to identify flooding events, we first compute the flow-duration exceedance probability ($\alpha$), which is the cumulative frequency curve that shows the probability of streamflow exceeding a given threshold for the data period [87]. In essence, this initial model predicts the probability that a flood of a certain magnitude occurs in the prediction window. We compute the annual flood exceedance probability using the historical streamflow data as:
$\alpha=m/(n+1)$ (5)
where, $m$ is the ranking from highest to lowest of annual maximum flow from the observed data, and $n$ is the total number of annual maximum streamflow cases. We identify the annual maximum flow with exceedance probability $20\%$, which is used as the flood threshold ($\gamma$). Therefore, any event with streamflow higher than the flood threshold is considered a flooding event.
3.4.1 Flood prediction
The flood risk indicator is a categorical value that classifies the likelihood of flooding based on the predicted streamflow values over the prediction horizon (n-steps) as shown in Figure 5. It is computed by comparing the predicted quantiles of streamflow (i.e. the expected value, $5^{th}$ percentile, and $95^{th}$ percentile predicted using the LSTM, Quantile-LSTMq=.05, Quantile-LSTMq=.95 models, respectively) against the predefined $\gamma$. For each of these models, we only consider the maximum predicted value over the prediction horizon to compute the flood risk indicator. The classification follows a three-tier system: (1) ’High Risk’ if the lower bound of the 90% confidence interval exceeds the threshold ($Q_{q=0.05}>\gamma$), (2) ’Moderate Risk’ if the expected value of streamflow surpasses the threshold ($Q_{reg}>\gamma$), and (3) ’Low Chance’ if only the upper bound ($Q_{q=0.05}>\gamma$) exceeds the threshold. If none of these conditions are met, the chance of flooding would be deemed ’unlikely’.
Figure 5: Ensemble model utilising Quantile-LSTM for prediction of flood probability and streamflow with uncertainty bounds using quantile regression.
@@ 3.5 Experiment Setup
In this section, we provide technical details for implementing the conventional deep learning models (LSTM, BD-LSTM, ED-LSTM, and CNN). Moreover, we also provide the details for our proposed methods, which include the flood risk indicator utilising Ensemble Quantile-LSTM. We determined model hyperparameters such as hidden neurons, learning rate and epochs from trial experiments, as shown in Table 2. We note that all the tested models implement a multi-step time series prediction with 5 inputs and 5 outputs (time steps). We train the model with a maximum time of 150 epochs using an Adam optimiser [88] with default hyperparameters, such as the initial learning rate.
We scaled the time-series data using the min-max-scaler, i.e. scaled in the range [0,1]. In each basin, there are about 10,000 data points representing daily data spanning from 1990 to 2020. We used the first 60 % of the data from 1990 as training data and the remaining for testing. We note that although there may be a climate bias in taking a block of data spanning more than a decade, this approach ensures reproducibility in results when compared to shuffling the data.
It is necessary to use a variety of performance metrics for model benchmarking, as no one evaluation metric can fully capture the consistency, reliability, accuracy, and precision of a stream flow model. The Nash–Sutcliffe-Efficiency (NSE) is the hydrologic analogue of the coefficient of determination ($R^{2}$ score), which can be less than zero. Silva et al. [89] have suggested the Squared Error Relevance (SER) - (SER-$\alpha$%) for extreme events. The SER calculates the Root Mean Squared Error (RMSE) of test windows that contain the $\alpha$ percentile of the data, measuring the ability of a model to predict rare data points. For example, SER-1% would find the RMSE of windows that contain the 1% most extreme stream flows for a given basin. Therefore, we use the following metrics to evaluate prediction performance accuracy over multiple steps (prediction horizon).
@@
Table 2: Configuration for conventional deep learning and conditional ensemble model.
Models Hidden Units Type
LSTM 20 Conventional Model
BD-LSTM 20 Conventional Model
ED-LSTM 20 Conventional Model
CNN 64 Conventional Model
Quantile-LSTMq=0.05 20 cells Quantile Ensemble
Quantile-LSTMq=0.95 20 cells Quantile ensemble
LSTM 20 cells Quantile Ensemble
Table 3: Accuracy (SER) on the test data of catchment strategies using LSTM model for South Australia
@@
Table 4: Architecture wise- model comparisons showing test accuracy (SER) to track best-performing model (South Australia)
@@
Quantile-LSTMq=0.05 2.0646 1.2959 0.6683 0.3790 0.1860 0.1109 0.0884 0.0801
±0.0580 ±0.0306 ±0.0154 ±0.0091 ±0.0045 ±0.0031 ±0.0028 ±0.0027
Quantile-LSTMq=0.95 1.5932 1.0813 0.6585 0.4390 0.2443 0.1478 0.1166 0.1051
±0.0459 ±0.0306 ±0.0169 ±0.0180 ±0.0114 ±0.0083 ±0.0080 ±0.0080
Table 5: Ensemble Quantile-LSTM model results showing test accuracy (SER) for major states in Australia
State Ensemble Quantile LSTM SER 1% SER 2% SER 5% SER 10% SER 25% SER 50% SER 75% RMSE
NT Regular LSTM 1.2810 0.8192 0.4845 0.3194 0.1621 0.0827 0.0618 0.0519
±0.0262 ±0.0190 ±0.0124 ±0.0078 ±0.0032 ±0.0016 ±0.0011 ±0.0008
Quantile LSTM q=0.05 1.8632 1.2321 0.7665 0.5174 0.2520 0.1275 0.0963 0.0816
±0.0078 ±0.0057 ±0.0041 ±0.0030 ±0.0016 ±0.0008 ±0.0005 ±0.0004
Quantile LSTM q=0.95 1.1136 0.7152 0.4377 0.3100 0.1810 0.0946 0.0705 0.0590
±0.0156 ±0.0097 ±0.0074 ±0.0092 ±0.0093 ±0.0053 ±0.0036 ±0.0028
NSW Regular LSTM 1.3857 0.9052 0.4724 0.2709 0.1214 0.0654 0.0449 0.0352
±0.0138 ±0.0100 ±0.0070 ±0.0050 ±0.0021 ±0.0010 ±0.0008 ±0.0006
Quantile LSTM q=0.05 1.5695 1.0295 0.5494 0.3163 0.1412 0.0752 0.0512 0.0398
±0.0144 ±0.0116 ±0.0085 ±0.0050 ±0.0020 ±0.0010 ±0.0007 ±0.0005
Quantile LSTM q=0.95 1.2753 0.8306 0.4350 0.2557 0.1194 0.0655 0.0452 0.0355
±0.0126 ±0.0085 ±0.0058 ±0.0054 ±0.0036 ±0.0020 ±0.0014 ±0.0011
ACT Regular LSTM 1.7718 1.1026 0.4953 0.2621 0.1182 0.0623 0.0423 0.0325
±0.0431 ±0.0230 ±0.0097 ±0.0055 ±0.0026 ±0.0015 ±0.0011 ±0.0009
Quantile LSTM q=0.05 1.9456 1.1943 0.5291 0.2756 0.1222 0.0637 0.0427 0.0326
±0.0221 ±0.0104 ±0.0059 ±0.0036 ±0.0017 ±0.0009 ±0.0006 ±0.0005
Quantile LSTM q=0.95 1.5312 1.0087 0.4852 0.2759 0.1355 0.0756 0.0529 0.0412
±0.0336 ±0.0231 ±0.0135 ±0.0101 ±0.0064 ±0.0042 ±0.0031 ±0.0025
WA Regular LSTM 2.2494 1.8241 1.1968 0.8022 0.4337 0.2719 0.2097 0.1809
±0.0519 ±0.0396 ±0.0229 ±0.0125 ±0.0034 ±0.0053 ±0.0039 ±0.0032
Quantile LSTM q=0.05 3.3430 2.7612 1.9196 1.2671 0.6159 0.3499 0.2675 0.2316
±0.0239 ±0.0200 ±0.0161 ±0.0126 ±0.0068 ±0.0038 ±0.0026 ±0.0019
Quantile LSTM q=0.95 1.7287 1.4083 0.9988 0.7465 0.5152 0.3706 0.2869 0.2423
±0.0327 ±0.0228 ±0.0152 ±0.0126 ±0.0166 ±0.0179 ±0.0148 ±0.0122
SA Regular LSTM 1.5887 1.0652 0.5973 0.3534 0.1778 0.0930 0.0633 0.0528
±0.0235 ±0.0158 ±0.0111 ±0.0084 ±0.0056 ±0.0031 ±0.0021 ±0.0018
Quantile LSTM q=0.05 2.2325 1.4956 0.8305 0.4787 0.2226 0.1151 0.0789 0.0674
±0.0124 ±0.0093 ±0.0062 ±0.0040 ±0.0018 ±0.0009 ±0.0006 ±0.0005
Quantile LSTM q=0.95 1.4409 1.0195 0.6268 0.4112 0.2318 0.1249 0.0854 0.0717
±0.0341 ±0.0265 ±0.0182 ±0.0144 ±0.0137 ±0.0081 ±0.0054 ±0.0043
QLD Regular LSTM 1.2907 0.8690 0.5061 0.3282 0.1559 0.0827 0.0557 0.0422
±0.0150 ±0.0110 ±0.0065 ±0.0041 ±0.0017 ±0.0008 ±0.0005 ±0.0004
Quantile LSTM q=0.05 1.6486 1.1355 0.6691 0.4291 0.2003 0.1053 0.0708 0.0536
±0.0183 ±0.0127 ±0.0088 ±0.0057 ±0.0026 ±0.0013 ±0.0009 ±0.0007
Quantile LSTM q=0.95 1.0673 0.7300 0.4512 0.3187 0.1709 0.0947 0.0645 0.0491
±0.0108 ±0.0051 ±0.0047 ±0.0050 ±0.0034 ±0.0019 ±0.0013 ±0.0010